OpenStudy (anonymous):

(y^2-1)(y^2+1)

OpenStudy (anonymous):

as (a^2 -B^2)=(a-b)(a+b)

OpenStudy (hoblos):

y^4-1 = (y^2-1)(y^2+1) = (y-1)(y+1)(y^2+1)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

can anyone explain please?

OpenStudy (anonymous):

i did so???

OpenStudy (hoblos):

y^4-1 = (y^2)^2 - 1^2 use the formula a^2 -B^2=(a-b)(a+b) so you get (y^2-1)(y^2+1) then use the formula again for (y^2-1) so the final answer would be (y-1)(y+1)(y^2+1)